# Adaptive dynamics for individual payoff game-theoretic models of vaccination

### v2

@@ 3 Solutions and equilibrium analysis
since $\pi(P)<1-P/p^{\ast}$ due to concavity of $\pi$. We have
@@ 4.1 Parameter $r$
    The only type of bifurcations that happens are pitchfork bifurcations, i.e. appearance of a degenerate equilibrium and its subsequent change into a stable-unstable-pair or a merge of a stable-unstable pair into a degenerate equilibrium and its disappearance.
@@ 4.2 Parameter $\epsilon$
Varying $\epsilon$ with results in pitchfork bifurcations of (3.1).
@@ 5 Examples
To this end, we change $g(x)$ accordingly: firstly, it needs to become a decreasing function. Secondly, we want it to be zero when $x\geq\frac{21}{22}$; thirdly, we want the measure of the locus of the points where it decreases to be small - say, $\frac{2}{11}$, in order to minimise the effect that the gluing function has on the probability of becoming infected.
To conform with these conditions, instead of $g(x)$ we must consider $\widetilde{g}(x):=1-g\left(11x-9\right)$. This function is 1 when $x\leq\frac{19}{22}$, monotonously decreases when $\frac{19}{22}\leq x\leq\frac{21}{22}$ and assumes zero value when $x>\frac{21}{22}$.
@@
This function is strictly monotonously decreasing and smooth. It is 1 when $p=1$ and turns and stays 0 when $p\geq\frac{4}{5}$. It’s second derivative changes sign, meaning that the function $\pi^{\prime}(P)$ is not monotonously increasing or decreasing.
Setting $r=\frac{1}{4}$ allows to plot the two functions, and it can be shown that they only intersect at one point, see Figure 4.1.
@@
$\begin{split}\pi^{\prime}(p)&=-(1-2x)^{2}\left(1-\int\limits_{-1}^{8(2x-1)}\exp\left(\frac{1}{t^{2}-1}\right)\mathrm{d}t\right)\\ &-\frac{16}{3}\exp\left(\frac{1}{(8(2x-1))^{2}-1}\right)\left((1-2x)^{3}+2\right)\end{split}$ (5.7)
@@
The former depicts the changes in the $r$-bifurcation diagram with varying values of $\epsilon$. The leftmost peak moves to the right with the increase in $\epsilon$ until it is no longer within the $[0,1]$ interval.
The second figure shows four $\epsilon$-bifurcation diagrams for various fixed values of $r$. Note that with the increase of $r$ the straight line curves, and the rightmost pitchfork point moves to the right until it disappears and the diagram becomes disjoint.

### v3

@@ 3 Solutions and equilibrium analysis
since $\pi(P)\leq 1-P/p^{\ast}$ due to concavity of $\pi$. We have
@@ 4.1 Parameter $r$
    The only type of bifurcations that happens is saddle-node, i.e. appearance of a degenerate equilibrium and its subsequent change into a stable-unstable-pair or a merge of a stable-unstable pair into a degenerate equilibrium and its disappearance.
@@ 4.2 Parameter $\epsilon$
Varying $\epsilon$ results in saddle-node bifurcations of (3.1).
@@ 5 Examples
To this end, we change $g(x)$ accordingly: firstly, it needs to become a decreasing function. Secondly, we want it to be zero when $x\geq\frac{4}{5}$; thirdly, we want the measure of the locus of the points where it decreases to be small - say, $\frac{1}{10}$, in order to minimise the effect that the gluing function has on the probability of becoming infected.
To conform with these conditions, instead of $g(x)$ we must consider $\widetilde{g}(x):=1-g\left(10x-5\right)$. This function is 1 when $x\leq\frac{7}{10}$, monotonously decreases when $\frac{7}{10}\leq x\leq\frac{4}{5}$ and assumes zero value when $x>\frac{4}{5}$.
@@
This function is strictly monotonously decreasing and smooth. It is 1 when $p=0$ and turns and stays 0 when $p\geq\frac{4}{5}$. Its second derivative changes sign, meaning that the function $\pi^{\prime}(P)$ is not monotonously increasing or decreasing.
Setting $r=\frac{1}{5}$ allows to plot the two functions, and it can be shown that they only intersect at one point, see Figure 4.1.
@@
$\begin{split}\pi^{\prime}(p)&=-2(1-2x)^{2}\left(1-\frac{\int\limits_{-1}^{8(2x-1)}\exp\left(\frac{1}{t^{2}-1}\right)\mathrm{d}t}{H(1)}\right)\\ &-\frac{16}{3}\exp\left(\frac{1}{(8(2x-1))^{2}-1}\right)\left((1-2x)^{3}+2\right)\frac{1}{H(1)}\end{split}$ (5.7)
Note that the number $\frac{1}{H(1)}$ is the coefficient from (5.3) and approximately equal to 2.2522836206907613.
@@
The former depicts the changes in the $r$-bifurcation diagram with varying values of $\epsilon$. The peak moves up with the increase in $\epsilon$ until it is no longer within the $[0,1]$ interval.
The second figure shows four $\epsilon$-bifurcation diagrams for various fixed values of $r$. Note that with the increase of $r$ the beginning of the curve becomes ”pinched”, and the new tip of the curve moves upwards until it disappears and the diagram becomes disjoint.